1a) It is desired to inflate a baggie with a volume of 936 milliliters by filling it with methane gas at a pressure of 1.18 atm and a temperature of 296 K. How many grams of CH4 gas are needed? _____ g

1b) A 0.705 gram sample of an unknown gas is found to occupy a volume of 0.620 L at a pressure of 472 mm Hg and a temperature of 59°C. The molecular weight of the unknown gas is ____ g/mol. Assume ideal behavior.

1c) A sample of hydrogen gas has a density of _____ g/L at a pressure of 0.890 atm and a temperature of 53 °C. Assume ideal behavior.
Answer #1

a) The ideal gas equation gives us the relation to find the number of moles of gas:

n = PV / RT, where P is pressure in atm, V is volume in Liters, R is gas constant and T Is temprature

Using the values given, we get: n = 1.18 * 0.936 / (0.0821 * 296)

n = 0.0455 moles

Mass of methane needed = moles of methane * molar mass = 0.0455 * 16 = 0.728 g

b) Using the ideal gas equation with P = 472 / 760 atm and T = 273 + 59 = 332 K

n = 472 * 0.620 / (760 * 0.0821 * 332) = 0.0141 moles

Number of moles = mass of gas / molar mass

Using this, molar mass = 0.705 / 0.0141 = 50 g/mol

c) Density of gas = PM / RT = 0.890 * 2 / (0.0821 * 326) = 0.067 g/L